OpenStudy (cagliero423):

Law of Sines help??? So I have a project and I have to find the measures of angles B and C on all these triangles. They give me measure A and what a and b equal and I know that sinA/a= SinB/b = SinC/c but every time I try and use this it doesn't work out, even when I have an example in my book and I type it into my calculator I get a different answer than the book and I'm just trying to understand what I'm doing wrong. Help would be appreciated!

OpenStudy (welshfella):

have you got an example?

OpenStudy (cagliero423):

Like in the book there is an equation and it says sin 109/RQ=sin39/40 RQ= 40 sin109/sin39 which is about 60.1 cm but when I type it into my calculator I get 33.9

OpenStudy (welshfella):

thats equation RQ = is correct

OpenStudy (cagliero423):

Yeah I know because it's in my book but I keep getting 33.9 and the answer is 60.1???

OpenStudy (cagliero423):

I type it in just as the equation says.

OpenStudy (welshfella):

ok calculators vary in the way you enter data with mine if in enter 40 * sin109 / sin 39 it gives me 60.09

OpenStudy (cagliero423):

that is seriously exactly what i type into mine? I don't understand hahah

OpenStudy (welshfella):

with some calculators you need to add angle first then sin

OpenStudy (welshfella):

try 109 then sine

OpenStudy (cagliero423):

Okay hold on

OpenStudy (cagliero423):

It says error

OpenStudy (welshfella):

ok the sin 109 is the way to go try sin 109 then hit exe

OpenStudy (cagliero423):

I don't see a button that says exe

OpenStudy (welshfella):

well whatever button you use to get a result

OpenStudy (cagliero423):

Okay I get .8167

OpenStudy (welshfella):

or does it give you the answer automatically

OpenStudy (welshfella):

well it should be 0.9455

OpenStudy (welshfella):

is your calculator set to degrees or radians?

OpenStudy (cagliero423):

I have no clue how do I find out???

OpenStudy (welshfella):

thats the problem!!!

OpenStudy (welshfella):

i just set mine to radians and got 0.816

OpenStudy (welshfella):

* 0.8167

OpenStudy (cagliero423):

omg, how do I change it to degrees?

OpenStudy (welshfella):

what make is ur calculator?

OpenStudy (cagliero423):

its a TI-84

OpenStudy (welshfella):

oh i don't know about those . Mines a Casio - maybe you can find a manual on-line.

OpenStudy (cagliero423):

Okay, I will do that, you're a life saver I would have never figured that out! haha thank you(:

OpenStudy (welshfella):

yw

OpenStudy (cagliero423):

Okay could you help me with something really quick actually? I just want to make sure I'm doing this right. my first equation I have angle A= 62 degrees a= 30 and b=32. So if I'm trying to find angle B I set it up like Sin62/30=SinB/32 right??

OpenStudy (cagliero423):

and then it's 32*Sin62/30 ?

OpenStudy (cagliero423):

@welshfella

OpenStudy (welshfella):

yes

OpenStudy (cagliero423):

Okay but then I get .9418 as my answer?

OpenStudy (welshfella):

By the way - i used to have a TI a long way back - i remember you changed degrees to radians and vise versa using the MODE key

OpenStudy (cagliero423):

Yeah I found a video on how to do it so i fixed that(:

OpenStudy (welshfella):

oh ok hold on

OpenStudy (cagliero423):

okie

OpenStudy (welshfella):

sin B = 0.9418 yes

OpenStudy (cagliero423):

Okay so then is angle B .9418 degrees? I don't get what to do next

OpenStudy (welshfella):

no thats the sine of the angle you need the inverse sine on ur calc - its either marked sin-1 or arcsin

OpenStudy (cagliero423):

Ohh okay, thank you again, I think that's all I needed to know(:

OpenStudy (welshfella):

yw